Evaluate limπ‘₯β†’2 π‘₯ 3βˆ’2π‘₯ 2+4π‘₯βˆ’8 π‘₯ 4βˆ’2π‘₯ 3+π‘₯βˆ’2


1
Expert's answer
2021-06-08T11:56:20-0400

Let us evaluate the limit:


"\lim\limits_{𝑥\to 2}\frac{ 𝑥^3−2𝑥^2+4𝑥−8}{ 𝑥^4−2𝑥^3+𝑥−2}=
 \lim\limits_{𝑥\to 2}\frac{ (x-2)(𝑥^2+4)}{ (x-2)(𝑥^3+1)}=
 \lim\limits_{𝑥\to 2}\frac{ 𝑥^2+4}{𝑥^3+1}=\frac{ 2^2+4}{2^3+1}=\frac{8}{9}."